Water flows through a pipe with an area of 13 in2 with a speed of 12 ft/s. It then enters a smaller pipe with an area of 3.1 in2. What is the speed of the water as it flows through the smaller pipe?
1
Expert's answer
2020-03-27T10:50:13-0400

Mechanics | Relativity

We need to find the speed of the water as it flows through the smaller pipe.

Solution:


Given,

Area of the first pipe = "13 \space in^2"


Speed of the first pipe = "12 \space ft/s"


Are of the small pipe = "3.1 \space in^2"


By conservation of mass and continuity equation


Area of first pipe "\times" velocity of the First Pipe = Area of small pipe "\times" velocity of small pipe


"13 \times 12 = 3.1 \times Velocity \space of \space small \space pipe"


velocity of small pipe

"= \frac {13 \times 12}{3.1} = \frac {156}{3.1} =50.32 \space ft/s"

Answer:

speed of the water flowing through the smaller pipe ="50.32 ft\/s"